# A Unified Multi-Wavelength Data Analysis Workflow with gammapy

### v1

@@ 4.2.2 Absorption components
[…]
  • Hydrogen I Absorption: Relevant for X-ray energies below $\sim 1\,\text{keV}$. Parameterized using the HI column density, following Wilms et al. (2000). For OP 313, the HI column density is approximately $1.25\times 10^{20}\,\text{cm}^{-2}$ (HI4PI Collaboration et al. 2016).
@@ 4.3 Forward folding fitting method
    Flux Point Fitting: This traditional method fits the model to observed flux points derived from analyzing each instrument’s data separately. Although it simplifies the fitting and can be computationally faster, it is observed to be more sensitive to the choice of initial parameters. Additionally, this approach may introduce errors due to differences in flux estimation and instrument calibration methodologies. It also overlooks upper limits and correlations between flux points, potentially affecting the accuracy of the fit.
@@
    Forward Folding Technique: This method fits the model directly to the counts or event data from the instruments, incorporating the IRFs in each iteration. It accounts for background components and full instrumental effects, leading to improved accuracy and, as observed, better convergence compared to the flux point fitting method.
@@
Other notable differences between the two methods include the treatment of instrumental effects and background components, more accurate in the forward folding case as the method directly incorporates energy redistribution and provides a more uniform representation of the emission in the observed counts space. In the NuSTAR case, bins for which the effective area of the instrument is approximately zero can still be of use as they can effectively constrain the instrumental background model. As opposed to directly fitting instrumental background components during the forward folding, the classical method inadequately performs a fit to ‘flux points’ implying that the background is either fixed or subtracted. A similar problem occurs with Fermi-LAT skymodel, consisting on the sum of diffuse components (isotropic component, galactic diffuse emission) and a number of point-like source lying close to OP 313. In the classical method, the contributions of such components are ‘frozen’, as opposed to fitted in the forward folding. Finally, the proposed workflow reduces systematic uncertainties associated with the conversion of counts to fluxes in the optical band as information coming from different filters are indirectly accounted for in the counts-flux conversion, likely improving the accuracy of the fit.
Figure 9: Joint multi-instrument forward-folding fit and joint fit using flux-points employing a phenomenological model with two distinct emission components and three absorption components. Top panel: individual datasets fitted with simple models (power-laws, log-parabolas, power-laws with exponential cutoffs) and corresponding reconstructed flux points, and joint fits (flux points, forward folding). Middle panel: Residuals, calculated as the ratio of reconstructed differential spectrum to model prediction for the flux points fit (Residuals PF, open circles), and as the significance or ratio of the difference between the excess counts and model predicted counts to the square root of the model predictions for each bin in the forward folding case (Residuals FF, small filled circles). For clarity, and due to the large amount of energy bins in the X-ray datasets, only one every 5 bins is shown for NuSTAR, and one every 3 bins is shown for XRT. Bottom panel: Multiplicative absorption components.
Figure 10: Corner plot showing the error ellipses calculated with a bootstrap method for the best-fit (free) parameters of the flux-point based fit and the full forward-folding joint fit. The model used for the LE is strongly degenerated because of the limited amount of available data to constrain that component.
@@ 5.1 Flexibility and modeling capabilities
This approach is particularly beneficial for X-ray data analysis, providing a modern and flexible framework compared to older tools like xspec, whose first release dates from 1983 and is often misused to produce flux points despite not being technically designed for that purpose. In the case of NuSTAR, this is especially relevant since instrumental background models quickly develop to a very complex sum of models components, making it almost impossible to extract flux points using xspec, as it would require to remove the contribution to the total flux from each of the components. The proposed flexible format, combined with tools such as gammapy, potentially enables a 3D analysis of XRT and NuSTAR data. This could be particularly advantageous for extended sources, or for a point source on top of an extended source (e.g. a halo), especially for instruments with limited angular resolution like NuSTAR. Therefore, the method tested in this work presents natural extensions and improvements.
@@ 5.2 Main challenges and practical limits
For Fermi-LAT, the current implementation assumes that the PSF depends only on energy, and is constant within the FoV. This is a potential limitation with respect to its native analysis pipeline, Fermitools. Fortunately, for bright point-like sources like OP 313, where the contribution of other sources in the FoV is almost negligible, this effect should be small. Being gammapy an active developed tool, this drawback may be addressed in the future.
@@ 5.3 Considerations for optical instruments
The approximate IRFs we introduced in this work still built and used in this study omit critical factors such as how atmospheric extinction, mirror reflectivity, aging of mirror coatings, and CCD defects modify the effective band profile. Including these elements in the IRFs — though technically challenging — would improve the accuracy of the spectral reconstruction. We did not explore spectral data analysis in the optical band due to the additional complexities involved, such as calibrating both wavelengths and fluxes, which could compromise the Poisson statistics of the observed data.
@@ 6 Conclusions
The resulting multi-instrument datasets and analysis code are stored in a format that is convenient for distribution and archiving. This format not only facilitates the reproduction of the results obtained in this specific study, but also supports further research on OP 313 with public datasets, as well as serve to make future validation tests for gammapy. Future work could involve detailed modeling of the spectral energy distribution using physical models through codes such as jetset (Tramacere et al. 2009, 2011), agnpy (Nigro et al. 2022), Bjet_MCMC (Hervet et al. 2024), or SOPRANO (Gasparyan et al. 2022, 2023).
This study provides a largely static view of OP 313 for two specific nights. Including variability studies and time-evolving models that account for particle injection and cooling will enhance our understanding of the emission of the source in particular and blazars in general. In this regard, the proposed simplified handling of instrument datasets will ensure greater consistency and reduce errors compared to classical methods that use data in disparate formats, as well as hopefully providing one of the first comprehensive multi-wavelength blazar datasets built over standarized data formats and open-source analysis tools.
@@ Acknowledgements.
We acknowledge the Fermi-LAT, NuSTAR, Swift-XRT, and Swift-UVOT teams for providing open-access data and analysis tools, as well as their support during this target-of-opportunity observation campaign. We also thank the Liverpool Telescope team for their prompt response and access to their robotic observations through reactive time. We are also grateful to the gammapy team for their extremely versatile and open-source analysis pipeline, which played a fundamental role in this multi-instrument, multi-wavelength project. M.N.R acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia, Innovación y Universidades (MCIU/AEI) under grant PARTICIPACIÓN DEL IAC EN EL EXPERIMENTO AMS and the European Regional Development Fund (ERDF) with reference PID2022-137810NB-C22 / DIO 10.13039/501100011033. J.O.S. acknowledges financial support through the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033 and through grants PID2019-107847RB-C44 and PID2022-139117NB-C44.
@@ Appendix A Software description and data repository
This work is based on the open-source analysis tool gammapy and open data formats initiatives (GADF, OGIP). Only two new code implementations were required to conduct the analysis as presented in this manuscript. The first is a reader for standard OGIP files generated by HEASARC’s native tools like uvot2pha, xrtproducts, and nuproducts. Originally developed by Giunti & Terrier (2022) for XMM, it was extended to also support NuSTAR, Swift-XRT, and Swift-UVOT. The second implementation is a set of Python classes that perform an optical point-like reduction photometry of Liverpool Telescope IO:O data, producing OGIP files from optical images using a set of reference stars with known photometry from Sloan’s SDSS DR18 and PanSTARR’s PS1 catalogs.
@@
Within the Notebooks/DatasetGenerator/ directory (with one subdirectory for each night of observations), we stored notebooks (one per instrument) that compile all the steps to generate the gammapy-compliant 1D and 3D binned datasets, as described in Section 3. Similarly, Notebooks/DatasetAnalysis/ contains notebooks with the analysis of the datasets from individual instruments and the joint analysis of the multi-wavelength (MWL) datasets. The Helpers/ directory contains functions for generating gammapy native multiplicative models for dust extinction, neutral hydrogen, and EBL absorption. It also includes classes for fetching photometry data from the PS1 and SDSS DR18 catalogs, as well as various file handling and plotting functions used across the analysis notebooks. The Models/ directory contains the tabular absorption models for EBL (Domínguez et al. 2011; Saldana-Lopez et al. 2021), the neutral hydrogen absorption exported from xspec’s TBabs model (NASA 2024a), and the dust extinction extracted from xspec’s redden (NASA 2024b). Finally, the Figures/ directory (one subdirectory per observing night) contains plots and figures presented in this work.
Appendix B Background estimation methods for NuSTAR
[…]
@@ Appendix C Dataset analysis validation
In addition, we show in Figure 11(b) the spectral reconstruction with both the instrumental background and the Poisson background estimate, well matching below $\sim 30$  keV. We note however that xspec has two severe limitations. The first is linked to the definition of a the power-law, whose pivot energy in xspec is fixed to 111\,1keV and cannot be changed. To avoid impacting the results with a reference energy that is outside the energy range of NuSTAR, we redefined the power-law so that the normalization is done with respect to the the integral flux in the $3-70\,$keV band, instead of at a specific energy. The second limitation is that xspec is not designed to produce flux points as we noted in Section 5.1, and only in simple cases with few spectral components (NuSTAR and Swift-XRT when doing an OFF background analysis with Poisson background statistics) it was feasible to estimate some sort of flux points for different energy bins. For more sophisticated cases, such as NuSTAR with the complex instrumental background model produced by nuskybgd, the flux points estimated with xspec do not correspond to the source component but to the total sky model, including the systematic background. Removing the additional background components (which are not fixed) is not trivial, therefore we only report in Figure 11(b) the SED confidence band from xspec, without the flux points. We also checked the parameter error contours in the amplitude vs spectral index parameter space in Figure 12, noting a very good agreement for the three instruments.
@@
Fermi-LAT NuSTAR A+B Swift-XRT
@111\,1GeV $3-70\,$ keV @111\,1keV
@@
native $1.00\pm 0.23$ $2.01\pm 0.16$ $6.30\pm 0.12$ $1.633\pm 0.030$ $9.28\pm 0.53$ $1.671\pm 0.075$
gammapy $0.98\pm 0.23$ $2.05\pm 0.19$ $6.33\pm 0.12$ $1.630\pm 0.030$ $9.09\pm 0.53$ $1.667\pm 0.077$
scale ×$\times 10^{-11}$$\times 10^{-4}$$\times 10^{-4}$
units $\mathrm{MeV^{-1}s^{-1}cm^{-2}}$ $\mathrm{s^{-1}cm^{-2}}$ $\mathrm{keV^{-1}s^{-1}cm^{-2}}$
@@
Figure 13: Best-fit spectra and SED points reconstructed using gammapy using counts and IRFs and with native tools using magnitudes or flux densities and multiplying by the effective filter frequency center. The emission in gammapy is modelled with a log-parabola with dust extinction. Flux points in gammapy are evaluated over the geometrical center of the bin (black filled circles) and at the effective mean energy of the band with open diamond symbols (under the assumption of a flat source photon spectrum, to be roughly comparable to the classical method). LT datasets from the same filter were stacked together using gammapy’s stack_reduce method, that co-adds the counts, background and exposures for datasets with compatible energy axes. For reference, weighted average fluxes for the native flux points are shown in colored red boxes, while brown error bars depict the total error once incorporating systematic uncertainties.
@@
Figure 14: Comparison of reconstructed ‘relative’ flux values using native photometry tools (which calculate source magnitudes or flux densities and multiply by the filter’s effective frequency) and those obtained with gammapy (using counts and instrument response functions). All ‘relative’ fluxes are shown as a ratio to the best-fit gammapy model of the joint datasets, evaluated at the corresponding point energy.
Appendix D Systematic Uncertainties in Liverpool Telescope Data
In classical analysis, systematic uncertainties related to errors in the zero-point estimation (star-to-star variations) are typically propagated in quadrature to the final error estimate of the source flux. While this method is model-independent, it does not take full advantage of the detailed broadband spectrum of the source, which becomes more apparent as additional data across different energy bands are incorporated.

### v2

@@ 4.2.2 Absorption components
  • ISM extinction: Relevant for optical and especially UV energies. Given OP 313’s Galactic latitude of $+83.34^{\circ}$ and a color excess of $E_{B-V}\approx 0.0125$ (Schlegel et al. 1998), interstellar medium (ISM) extinction is minimal below $10$ eV. At $\sim 5\,\text{eV}$ (UVOT’s UVM2 filter), the transparency is approximately 90% (Cardelli et al. 1989).
  • Hydrogen I absorption: Relevant for X-ray energies below $\sim 1\,\text{keV}$. Parameterized using the neutral hydrogen (HI) column density Wilms et al. (2000), which is $1.25\times 10^{20}\,\text{cm}^{-2}$ for OP 313 (HI4PI Collaboration et al. 2016).
@@ 4.3 Forward folding fitting method
    Flux Point Fitting: This traditional method fits the model to observed flux points derived from analyzing each instrument’s data separately. Although it simplifies the fitting and can be computationally faster, it is observed to be more sensitive to the choice of initial parameters. Additionally, this approach may introduce errors due to differences in flux estimation and instrument calibration methodologies. It also overlooks upper limits and correlations between flux points.
@@
    Forward Folding Technique: This method fits the model directly to the counts or event data from the instruments, convolving with the IRFs in each iteration. It allows incorporating background components and full instrumental effects.
@@
Other notable differences between the two methods include the treatment of instrumental effects and background components, more accurate in the forward folding case as the method directly incorporates energy redistribution and provides a more uniform representation of the emission in the observed counts space. In the NuSTAR case, bins for which the effective area of the instrument is approximately zero can still provide useful constraints to the instrumental background model. As opposed to directly fitting instrumental background components during the forward folding, the classical method inadequately performs a fit to ‘flux points’ implying that the background is either fixed or subtracted. A similar problem occurs with Fermi-LAT skymodel, consisting on the sum of diffuse components (isotropic component, galactic diffuse emission) and a number of point-like source lying close to OP 313. In the classical method, the contributions of such components are ‘frozen’, as opposed to fitted in the forward folding. Finally, the proposed workflow reduces systematic uncertainties associated with the conversion of counts to fluxes in the optical band as information coming from different filters is indirectly accounted for in the counts-flux conversion, likely improving the accuracy of the fit.
Figure 9: Joint multi-instrument forward-folding fit and joint fit using flux-points employing a phenomenological model with two distinct emission components and three absorption components. Top panel: individual datasets fitted with simple models (power-laws, log-parabolas, power-laws with exponential cutoffs) and corresponding reconstructed flux points, and joint fits (flux points, forward folding). Middle panel: Residuals, calculated as the ratio of reconstructed differential spectrum to model prediction for the flux points fit (Residuals PF, open markers), and as the significance or ratio of the difference between the excess counts and model predicted counts to the square root of the model predictions for each bin in the forward folding case (Residuals FF, small filled circles). For clarity, and due to the large amount of energy bins in the X-ray datasets, only one every 5 bins is shown for NuSTAR, and one every 3 bins is shown for XRT. Bottom panel: Multiplicative absorption components.
Figure 10: Corner plot showing the error ellipses calculated with a bootstrap method for the best-fit (free) parameters of the flux-point based fit and the full forward-folding joint fit. The model used for the LE is highly degenerate because of the limited amount of available data to constrain that component.
@@ 5.1 Flexibility and modeling capabilities
The approach proposed here for multi-wavelength fitting is particularly beneficial for the X-ray data analysis part in this broadband context. Classical tools like xspec are unfortunately often misused in the $\gamma$-ray community to produce flux points that are later fitted in the multi-wavelength context (e.g. Nigro et al. 2022; Alfaro et al. 2022) despite not being technically designed for that purpose. In the case of NuSTAR, this is especially relevant since its instrumental background model is a complex sum of multiple components, making it almost impossible to extract for the source of interest flux points using xspec, as it would require to remove the contribution to the total flux from each of the other components. The proposed flexible framework, based on gammapy, removes the need of multi-wavelength flux points, and potentially enables a 3D analysis of XRT and NuSTAR data. The latter could be particularly advantageous for extended sources, or for a point source on top of an extended source (e.g. a halo), especially for instruments with limited angular resolution like NuSTAR. Therefore, the method tested in this work presents natural extensions and improvements.
@@ 5.2 Main challenges and practical limits
For Fermi-LAT, our analysis assumes that the PSF depends only on energy, and is constant within the FoV. This is a potential limitation with respect to its native analysis pipeline, Fermitools. Fortunately, for bright point-like sources like OP 313, where the contribution of other sources in the FoV is almost negligible, this effect should be small. Being gammapy an active developed tool, this drawback may be addressed in the future.
As photon-counting devices, Swift-XRT and Swift-UVOT are subject to the effects of pile-up and coincidence loss, which happen when the photon flux is large enough that two or more photons have a significant probability of being detected at the same position within a single CCD readout frame. In such cases, the detector registers the multiple events as a single event, with the recorded energy being the sum of the overlapping photons. This phenomenon skews the observed energy distribution and modifies the statistical properties of the data, deviating from pure Poissonian behavior and affecting the accuracy of count rate uncertainties.
The treatment of these effects varies per instrument. For UVOT, a correction for coincidence loss is implemented in uvotsource, effective as long as the field is not crowded and the source remains point-like (Breeveld et al. 2010). In the case of XRT, pile-up is less common in Windowed Timing (WT) mode unless the source is very bright, though it requires specialized treatment when it does occur (Mineo et al. 2006). In Photon Counting (PC) mode (Vaughan et al. 2006), a common approach to mitigate pile-up involves defining an annular extraction region to mask the brightest central pixels and estimate the source counts from the wings of the PSF. Like in UVOT, this approach is feasible only for isolated, point-like sources with minimal background contamination. More complex scenarios, such as extended sources or crowded fields, would likely require a more sophisticated spatial and spectral modeling of the emission and pile-up effects, well beyond the scope of this work.
@@ 5.3 Considerations for optical instruments
The approximate IRFs we introduced in this work omit critical factors such as how atmospheric extinction, mirror reflectivity, aging of mirror coatings, and CCD defects modify the effective band profile. Including these elements in the IRFs — though technically challenging — would improve the accuracy of the spectral reconstruction. We did not explore spectral data analysis in the optical band due to the additional complexities involved, such as calibrating both wavelengths and fluxes, which could compromise the Poisson statistics of the observed data.
@@ 6 Conclusions
The resulting multi-instrument datasets and analysis code are stored in a format that is convenient for distribution and archiving. This format not only facilitates the reproduction of the results obtained in this specific study, but also supports further research on OP 313 with public datasets, as well as serve to make future validation tests for gammapy. Future work could involve detailed modeling of the spectral energy distribution using physical models through codes such as jetset (Tramacere et al. 2009, 2011), agnpy (Nigro et al. 2022), Bjet_MCMC (Hervet et al. 2024; Hervet 2024), or SOPRANO (Gasparyan et al. 2022, 2023), or exploring nested sampling methods (Buchner 2021).
This study provides a largely static view of OP 313 for two specific nights. Including variability studies and time-evolving models that account for particle injection and cooling will enhance our understanding of the emission of the source in particular and blazars in general. In this regard, the proposed simplified handling of instrument datasets will ensure greater consistency and reduce errors compared to classical methods that use data in disparate formats, as well as hopefully providing one of the first comprehensive multi-wavelength blazar datasets built over standardized data formats and open-source analysis tools.
@@ Acknowledgements.
We acknowledge the Fermi-LAT, NuSTAR, Swift-XRT, and Swift-UVOT teams for providing open-access data and analysis tools, as well as their support during this target-of-opportunity observation campaign. We also thank the Liverpool Telescope team for their prompt response and access to their robotic observations through reactive time. We are also grateful to the gammapy team for their extremely versatile and open-source analysis pipeline, which played a fundamental role in this multi-instrument, multi-wavelength project. M.N.R. acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia, Innovación y Universidades (MCIU/AEI) under grant PARTICIPACIÓN DEL IAC EN EL EXPERIMENTO AMS and the European Regional Development Fund (ERDF) with reference PID2022-137810NB-C22 / DIO 10.13039/501100011033. J.O.S. and D.M. acknowledge financial support from the projetct ref. AST22_00001_9 with founding from the European Union - NextGenerationEU, the Ministerio de Ciencia, Innovación y Universidades, Plan de Recuperación, Transformación y Resiliencia, the Consejería de Universidad, Investigación e Innovación from the Junta de Andalucía and the Consejo Superior de Investigaciones Científicas. J.O.S. also acknowledges financial support through the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033 and through grants PID2019-107847RB-C44 and PID2022-139117NB-C44 and founding from INFN Cap. U.1.01.01.01.009. F.A. acknowledge support from the European Union’s Horizon 2020 Programme under the AHEAD2020 project (grant agreement n. 871158) and support by CNES, focused on methodology for X-ray analysis.
@@ Appendix A Software description and data repository
This work is based on the open-source analysis tool gammapy and open data formats initiatives (GADF, OGIP). Only two new code implementations were required to conduct the analysis as presented in this manuscript. The first is a reader for standard OGIP files generated by HEASARC’s native tools like uvot2pha, xrtproducts, and nuproducts. This reader was originally developed by Giunti & Terrier (2022) for XMM, and it was extended to also support NuSTAR, Swift-XRT, and Swift-UVOT. The second implementation is a set of Python classes that perform an optical point-like photometric reduction of Liverpool Telescope IO:O data, producing OGIP-compliant files from optical images using a set of reference stars with known photometry from Sloan’s SDSS DR18 and PanSTARR’s PS1 catalogs.
@@
Within the Notebooks/DatasetGenerator/ directory (with one subdirectory for each night of observations), we stored notebooks (one per instrument) that compile all the steps to generate the gammapy-compliant 1D and 3D binned datasets, as described in section 3. Similarly, Notebooks/DatasetAnalysis/ contains notebooks with the analysis of the datasets from individual instruments and the joint analysis of the multi-wavelength (MWL) datasets. The Helpers/ directory contains functions for generating gammapy native multiplicative models for dust extinction, neutral hydrogen, and EBL absorption. It also includes classes for fetching photometry data from the PS1 and SDSS DR18 catalogs, as well as various file handling and plotting functions used across the analysis notebooks. The Models/ directory contains the tabular absorption models for EBL (Domínguez et al. 2011; Saldana-Lopez et al. 2021), the neutral hydrogen absorption exported from xspec’s TBabs model (Wilms et al. 2000; NASA 2024), and the dust extinction extracted from xspec’s redden (Cardelli et al. 1989; NASA 2024). Finally, the Figures/ directory (one subdirectory per observing night) contains plots and figures presented in this work.
Appendix B Reproducibility and Analysis Tools
To ensure reproducibility of the multi-wavelength analysis presented in this work, detailed information is provided on the software tools (including versions), calibration files, observation IDs used for each instrument. The data were processed using publicly available software and calibrated with the most up-to-date instrument response files. The following tables summarize the relevant details for each instrument.
B.1 Fermi-LAT
Fermi-LAT data were analyzed using the Fermitools package version 2.2.0, using the latest available diffuse models for PASS8 and the most recent 4FGL-DR4 catalog release at the time. The analysis settings are shown in table 1.
Table 1: Details for Fermi-LAT data analysis.
Parameter Details
Software Fermitools 2.2.0 (Conda), enrico (git, April 2024)
Data release PASS-8
Diffuse Model gll_iem_v07, iso_P8R3_SOURCE_V3_v1
Catalog (skymodel) 4FGL-DR4 (gll_psc_v35)
Sources (skymodel) Those within $<25^{\circ}$ with Test Statistics $TS>1$, free spectral parameters if $<5^{\circ}$
B.2 NuSTAR
NuSTAR observations were analyzed using the nustardas package version 2.1.2, as part of the HEASoft 6.31.1 software suite. The instrumental background modeling was performed with nuskybgd. The observation IDs and analysis details are listed in table 2
Table 2: Details for NuSTAR data analysis.
Parameter Details
Software HEASoft V6.31.1, nustardas 2.1.2, nuproducts 0.3.3
Background model nuskybgd 0.1.dev206
CALDB 20240729
Obs. IDs 91002609002 (MJD 60373), 91002609004 (MJD 60384)
B.3 Swift-XRT
The X-ray Telescope (XRT) data were analyzed using the tools provided in the HEASoft package, version 6.32.1. Specific details on the software versions, observations, and response files are reported in table 3.
Table 3: Details for Swift-XRT data analysis.
Parameter Details
Software HEASoft 6.32.1, xrtproducts 0.4.3, XSPEC 12.13.1
CALDB 20240729
Response Matrices swxpc0to12s6_20210101v016
Obs. IDs (OBSID) 00036384074 (MJD 60373), 00089816002 (MJD 60384)
B.4 Swift-UVOT
The UVOT data were processed using the UVOT software suite within HEASoft 6.32.1, and the final data products were extracted using uvot2pha, which internally uses uvotproduct and uvotsource. Specific details are summarized in Table 4.
Table 4: Details for Swift-UVOT data analysis.
Parameter Details
Software HEASoft 6.32.1, uvotproduct 2.9, uvotsource 4.5, uvot2pha 1.1
CALDB 20240729
Response File (RSP) 20041120v105
Obs. IDs 00036384074 (MJD 60373), 00089816002 (MJD 60384)
B.5 Liverpool IO:O
The Liverpool IO:O was provided in already reduced format by Liverpool’s automatic reduction, and the final photometric analysis was performed with a custom-made code, provided in the Notebooks folder of the GitHub repository, which is based on common python packages like astropy or photutils, and public catalogs from Sloan and Pan-STARRS
Table 5: Details for Liverpool IO:O data analysis.
Parameter Details
Software photutils 1.8.0, astropy 5.3.4, astroquery 0.4.7, gammapy 1.3.dev726+g6c6d15956
Catalogs Pan-STARRS PS1-DR2, SDSS DR18
@@ Appendix C Dataset analysis validation
In addition, we show in Figure 11(b) the spectral reconstruction with both the instrumental background and the Poisson background estimate, matching well below $\sim 30$  keV. We note however that xspec has two severe limitations. The first is linked to the definition of a the power-law, whose pivot energy in xspec is fixed to 111\,1keV and cannot be changed. To avoid impacting the results with a reference energy that is outside the energy range of NuSTAR, we redefined the power-law so that the normalization is done with respect to the the integral flux in the $3-70\,$keV band, instead of at a specific energy. The second limitation is that xspec is not designed to produce flux points as we noted in Section 5.1, and only in simple cases with few spectral components (NuSTAR and Swift-XRT when doing an OFF background analysis with Poisson background statistics) it was feasible to estimate some sort of flux points for different energy bins. For more sophisticated cases, such as NuSTAR with the complex instrumental background model produced by nuskybgd, the flux points estimated with xspec do not correspond to the source component but to the total sky model, including the systematic background. Removing the additional background components (which are not fixed) is not trivial, therefore we only report in Figure 11(b) the SED confidence band from xspec, without the flux points. We also checked the parameter error contours in the amplitude vs spectral index parameter space in Figure 12, noting a very good agreement for the three instruments.
@@
Fermi-LAT (at 111\,1GeV) NuSTAR A+B ($3-70\,$ keV) Swift-XRT (at 111\,1keV)
@@
native $1.00\pm 0.23$ $2.01\pm 0.16$ $6.30\pm 0.12$ $1.633\pm 0.030$ $10.20\pm 0.60$ $1.683\pm 0.077$
gammapy $0.98\pm 0.23$ $2.05\pm 0.19$ $6.32\pm 0.12$ $1.630\pm 0.030$ $10.29\pm 0.60$ $1.690\pm 0.078$
scale/units ×$\times 10^{-11}$ $\mathrm{MeV^{-1}s^{-1}cm^{-2}}$$\times 10^{-4}$ $\mathrm{s^{-1}cm^{-2}}$$\times 10^{-4}$ $\mathrm{keV^{-1}s^{-1}cm^{-2}}$
@@
Finally, to validate the analysis on single-channel datasets, we estimated the all-channel best-fit spectral model for both Swift-UVOT and Liverpool IO:O, together with reconstructed flux points for each filter individually. The latter are compared to the flux derived from the native analysis, i.e. the output of uvotsource for Swift-UVOT, and the differential photometry fluxes reconstructed with our photutils-based pipeline for Liverpool IO:O. The resulting SEDs are shown in figures 13(a) and 13(b) respectively. To reduce the effect introduced by the ambiguity in the selection of the reference energy of the band (estimated differently for gammapy and the native analyses), figures 14(a) and 14(b) compare the ratio of each flux point over the best model prediction at the reference energy of the flux point.
@@
Figure 13: Best-fit spectra (log-parabola with dust extinction) and flux points reconstructed using gammapy — based on counts and IRFs — and with native tools using flux densities multiplied by the effective filter frequency center. Flux points in gammapy are evaluated over the geometrical center of the bin (black filled circles) and at the effective mean energy of the band with open diamond symbols (under the assumption of a flat source photon spectrum, to be roughly comparable to the classical method). LT datasets from the same filter were stacked together using gammapy’s stack_reduce method, that co-adds the counts, background and exposures for datasets with compatible energy axes. For the native analysis, weighted averages are shown instead as red boxes. Dashed gray error bars depict the total error once incorporating systematic uncertainties.
@@
Figure 14: Comparison of reconstructed ‘relative’ flux values using native photometry tools (reconstructed flux densities multiplied by the filter’s effective frequency) and those obtained with gammapy (using counts and instrument response functions). All ‘relative’ fluxes are shown as a ratio to the best-fit gammapy model of the joint datasets, evaluated at the corresponding point energy.
Appendix D Background estimation for NuSTAR
[…]
Appendix E Systematic Uncertainties in Liverpool Telescope Data
In the classical analysis, systematic uncertainties related to errors in the zero-point estimation (star-to-star variations) are typically propagated in quadrature to the final error estimate of the source flux. While this method is model-independent, it does not take full advantage of the detailed broadband spectrum of the source, which becomes more apparent as additional data across different energy bands are incorporated.